OpenStudy (anonymous):

I need help ASAP! I will fan and medal anyone who helps me as QUICKLY as possible. 1. Given F(x) = 3x 2 - 4, find F(-√3). 2. Given F(x) = 3x 2 - 4 and F(a) = 8, find a. 3. If F(x) = 2x and G(x) = log x, find F [G(x)]. 4. If F(x) = 3 log x and G(x) = x 2 + 1, find G[F(x)]. 5. If F(x) = log 5 2x, find F -1(x). 6. Which of the following is a rational zero of the function P(x) = 4x 4 + 8x 3 - 7x 2 - 21x - 9? A. -3/2 B. -1 C. -3/4

OpenStudy (jadeishere):

I'm not sure how to help you past the first one

OpenStudy (31356):

On the first one is it 3x+2-4 or 3x^2 -4?

OpenStudy (jadeishere):

it's 3x^2

OpenStudy (anonymous):

Yeah, Its raised 2.

OpenStudy (31356):

Oh okay....

OpenStudy (anonymous):

I just have no clue how to do any of this :c

OpenStudy (ipwnbunnies):

For the first question, plug in -sqrt(3) for x.

OpenStudy (ipwnbunnies):

\[3(-\sqrt{3})^2 - 4 = ?\]

OpenStudy (anonymous):

I got the first one, it was close to five, and that was one of my answers so I chose it haha

OpenStudy (ipwnbunnies):

Oh, for the second question, set the function equal to 8. 8 = 3x^2 - 4 Can you solve for x?

OpenStudy (ipwnbunnies):

Err, which ones are we doing then lol.

OpenStudy (anonymous):

All of them, besides the first one haha

OpenStudy (ipwnbunnies):

Alright, do the second one then lol. It's ez peazy.

OpenStudy (anonymous):

I am dumbbb. I dont know how to get x because of the squared

OpenStudy (ipwnbunnies):

Can you do something else first? With the 4?

OpenStudy (anonymous):

Add it to the 8 so it is 12=3x^2

OpenStudy (ipwnbunnies):

Good, then, divide right?

OpenStudy (anonymous):

by three and get 4=x^2, the square root and get 2?

OpenStudy (ipwnbunnies):

Not just 2 buttttt

OpenStudy (anonymous):

x=2 ?

OpenStudy (ipwnbunnies):

When you square root something, you will get the positive of that, like 2. But you can also get the negative of that, -2. (2)^2 = 4 (-2)^2 = 4

OpenStudy (ipwnbunnies):

The question didn't ask for all values of a, so it's w/e lol

OpenStudy (anonymous):

Ohh!!! Okay! That was one of my answers so, next question! :D

OpenStudy (ipwnbunnies):

f(g(x)) means we're putting the function g(x) into the x's of f(x). :3 Get my drift?

OpenStudy (anonymous):

Yes, kinda :P

OpenStudy (ipwnbunnies):

Let's see if you can do that one then. Show me what you got.

OpenStudy (anonymous):

Uhmm, 2logx?

OpenStudy (ipwnbunnies):

#4 is also the same deal.

OpenStudy (ipwnbunnies):

Correct :)

OpenStudy (anonymous):

okay :D We're on a roll! haha

OpenStudy (anonymous):

I think i got the rest of them! Thank you soo much!! you have saved me!!!! <3 <3 <3